Determine the limit when x goes to zero (x^(2)+5x-4)/(x^(2)+1) if it exists.


1
Expert's answer
2021-01-29T12:41:05-0500

"\lim \limits_{x \to 0} \frac{x^{2} + 5x - 4}{x^{2} + 1} = \frac {(0)^{2} + 5(0) - 4}{(0)^{2} + 1} = \frac{0+0-4}{0+1}"

"\lim \limits_{x \to 0} \frac{x^{2} + 5x - 4}{x^{2} + 1} = \frac{-4}{1} = -4"